OpenStudy (anonymous):

Could someone helo me? I'm learning about quadratic equations, and I'm not quite sure how to set this problem up to solve. A= pi r^2 + pi rs

OpenStudy (anonymous):

I have to solve for r.

OpenStudy (anonymous):

nm, I thought it was set equal to zero.. Do you know quadratic formula?

OpenStudy (anonymous):

You can rewrite it in the form, ax^2 +bx +c : a=π, b=sπ, c=-A

OpenStudy (anonymous):

okay, I knew i had to plug in a + b = c , not unsure of the direction I needed to go

OpenStudy (anonymous):

so would it look like this ? ax^2 + bx + c = pi(x)^2 + (pi)sx +(pi)-x^2 ?

OpenStudy (anonymous):

Almost. \[\pi r^2 +s \pi r -A=0.\]

OpenStudy (anonymous):

'A' is different from 'a' here.

OpenStudy (anonymous):

how do I show the pi symbol when posting on here?

OpenStudy (anonymous):

because A is area right?

OpenStudy (anonymous):

If you're using a PC, hold down ALT key, press 227 on the numeric keypad then release ALT.

OpenStudy (anonymous):

hmm so if I'm using a laptop it's not possible?

OpenStudy (anonymous):

Yes, assuming this is some sort of area formula, I'd guess A meant area. There's no context, and it doesn't really matter. My laptop has a numpad, others that don't usually have some sort of function key that lets you use other keys for the numpad.

OpenStudy (anonymous):

If you're using a Macbook or similar, they have their own peculiar codes for symbols. I don't know any of them, but you can look them up online. Otherwise, you can open the Equation editor here and just type pi and it'll do it.

OpenStudy (anonymous):

oh okay, thanks for the tip